Solve by the substitution method. What is the solution to the system?
x-y=3
7x+3y=-19

Answer by Boreal(15235)

the top is also x=y+3
so the second is 7(y+3)+3y=-19
7y+21+3y=-19
10y=-40
y=-4
x=-1
-
check in second -7-12=-19